# Doppler Pulse Amplification

### v1

@@ Abstract
The ability to amplify ultrashort pulses has revolutionized modern laser science, driving advances in various fields such as ultrafast optics and spectroscopy. A pivotal development in this field is chirped pulse amplification (CPA), which stretches, amplifies and recompresses ultrashort optical pulses using dispersive elements to overcome amplification limits. However, CPA faces limitations due to gain narrowing, restricting the final pulse duration. Here, we propose Doppler pulse amplification (DoPA), a novel approach for amplifying ultrashort pulses. While DoPA shares similarities with CPA in that it also stretches, amplifies and recompresses pulses, it differs in how it achieves temporal compansion. Unlike CPA, DoPA exploits Doppler shifts induced by space-time modulated interfaces through a space-time wedge implementation without chirping. We show that DoPA dynamically shifts the pulse spectrum, effectively mitigating the gain narrowing issue of CPA. Additionally, we show that DoPA enables more compact amplification systems via a space-time Fresnel implementation, enabling amplification of picosecond and femtosecond pulses. This approach may pave the way for more efficient, high-intensity laser systems and expand the potential for applications in both laboratory research and practical environments.
@@ I Introduction
Pulse amplification techniques have continuously evolved to meet the growing demands of various fields, from telecommunications and medical applications to scientific research. Some applications, such as communication [1, 2] and radar systems [3, 4], require only moderate gain amplification. However, others—including laser light-matter interactions [5, 6], particle acceleration [7, 8], ultrafast optics [6, 9] and precision eye surgery [10, 11]—demand advanced techniques to amplify ultrashort pulses to very high peak powers. In these cases, directly amplifying ultrashort pulses exceed the saturation level of the amplifier, potentially damaging optical components and introducing nonlinear effects such as self-phase modulation [12, 13]. To address this, two main approaches have been developed. The first, chirped pulse amplification (CPA) [14, 15, 16, 17, 18], uses dispersive elements, such as diffraction gratings, to temporally stretch the pulse, reducing its peak power and spreading its energy over time. This enables greater amplification before the pulse is recompressed to its original shape and duration, now with significantly higher peak power. Developed by Strickland and Mourou, CPA revolutionized laser science and earned them the 2018 Nobel Prize in Physics [14, 15]. The second approach, divided-pulse amplification (DPA) [19, 20, 21, 22, 23, 24, 25], splits the incident pulse into multiple identical lower-power copies using a sequence of birefringent crystals. These copies are individually amplified and then recombined, resulting in a pulse with significantly higher peak power.
@@
Despite their success, both approaches come with intrinsic limitations that restrict their amplification performance. On one hand, CPA suffers from gain narrowing, a spectral bottleneck that arises when the spectral width of the pulse exceeds the bandwidth of the gain amplifier, leading to temporal broadening and pulse distortion in the output [26, 27, 18]. Additionally, CPA is constrained by diffraction grating sizes, making it impractical for pulses longer than femtoseconds [28, 21]. On the other hand, DPA is limited by the number of pulse divisions, restricting the reduction in peak power of the individual copies. Consequently, amplification is constrained, as the copies quickly reaches the saturation level of the amplifier [29, 30, 31].
@@
We propose a novel method for amplifying ultrashort pulses, called Doppler pulse amplification (DoPA). DoPA shares the same top-level structure with CPA—stretching, amplifying and compressing—but instead of using dispersive elements, it relies on Doppler shifts[32] induced by moving modulated interfaces [33, 34, 35, 36] to control the pulse compansion, without any chirping. It mitigates the gain narrowing problem by dynamically alternating the pulse spectrum and may lead to compact systems, making it suitable for the amplification of both picosecond and femtosecond pulses.
@@
The paper is structured as follows. Section II introduces DoPA in parallel to CPA. Section III analyzes the DoPA system using a space-time wedge implementation. Section IV demonstrates this analysis. Section V explores how the device size can be reduced by replacing the space-time wedge with a space-time Fresnel profile. Finally, Sec. VI concludes the paper and discusses potential experimental implementations.
@@
Figure 1: General concept of (a) chirped pulse amplification (CPA) and (b) proposed Doppler pulse amplification (DoPA). In both CPA and DoPA, the input pulse is first stretched out in time to reduce its peak power from $P_{\text{in}}$ to $P^{\prime}_{\text{in}}=P_{\text{in}}/a$, where $a>1$ is the expansion factor. This temporal stretching—achieved via a chirp filter in CPA and an opening space-time wedge in DoPA—lowers the peak power enough to allow amplification up to the saturation threshold, $P_{\text{th}}$, of the amplifier. With gain, $G_{\text{a}}$, the amplified stretched pulse reaches $P^{\prime}_{\text{out}}=G_{\text{a}}P^{\prime}_{\text{in}}=P_{\text{th}}$. The pulse is then recompressed using a matched chirp filter (CPA) and a closing space-time wedge (DoPA), restoring its duration and boosting the peak power to $P_{\text{out}}=aP^{\prime}_{\text{out}}=G_{\text{a}}P_{\text{in}}$. Unlike CPA, which requires a broad amplifier bandwidth, DoPA compresses the spectrum during the stretching stage, enabling amplification with a much narrower-band amplifier.
@@ II Concept
Although based on a fundamentally different mechanism, DoPA shares the principle of temporal compansion (expansion and compression) with CPA. Therefore, it makes sense to introduce DoPA alongside CPA. Figure 1 depicts the two approaches, with Fig. 1a and 1b corresponding to CPA and DoPA, respectively.
@@
A CPA system, shown in Fig. 1a, consists of three main stages [14]. First, the input pulse—with peak power $P_{\text{in}}$—passes through a chirp filter characterized by a phase modulation $\varphi\left[t\right]$, whose group velocity dispersion expands and chirps the pulse [9]. This causes different frequency components to travel at different speeds, resulting in temporal stretching of the pulse. This chirping process reduces the peak power from $P_{\text{in}}$ to $P_{\text{in}}/a$, where $a>1$ is the compansion factor while spreading its energy over a longer duration. Secondly, the stretched pulse is amplified by a gain amplifier, e.g., Ti:sapphire or Nd:glass, up to its saturation threshold, $P_{\text{th}}$, to mitigate nonlinear effects and prevent damage to the optical components, resulting in a peak power $P^{\prime}_{\text{out}}=G_{\text{a}}P^{\prime}_{\text{in}}$, where $G_{\text{a}}$ is the amplifier gain. Finally, a second chirp filter, with opposite dispersion characteristics, compresses the pulse back to its original duration, eliminating the chirp introduced in the first step and significantly increasing the output peak power to $P_{\text{out}}=G_{\text{a}}P_{\text{in}}$.
@@
Since the incident pulse is ultrashort, it possesses a broad spectral width. In the case of CPA, this width remains unchanged by the chirp filter—only its shape is altered [9]. This imposes strict constraints on the amplifier, whose limited bandwidth may result in distortions, as only a portion of the spectrum is amplified, undermining the original goal of the system. This phenomenon, known as gain narrowing, is one of the principal limitations of CPA. In the figure, it is avoided by employing a very broadband amplifier, but it would arise if the narrower-band amplifier corresponding to the dotted curve were used [26, 27, 18].
@@
While DoPA also involves sequential pulse expansion, amplification and recompression, as illustrated in Fig. 1b, it achieves these transformations through Doppler frequency shifts induced by moving interfaces, instead of chirp filters. DoPA employs two paired space-time wedges, recently introduced in [36]. The wedge edges are assumed to act as perfect electric conductor moving interfaces, ensuring that the pulse remains confined within the structure. This configuration allows the pulse to bounce back and forth between two moving interfaces and depending on the configuration—–either opening or closing—–the pulse either expands or compresses upon interaction with the interfaces. In the case of an opening wedge, the pulse interacts only with comoving interfaces, leading to Doppler downshifting, which stretches the pulse in time by a factor $a$ and reduces its peak power to $P_{\text{in}}/a$, facilitating safe amplification. Conversely, for a closing wedge, the wave encounters only contramoving interfaces, resulting in Doppler upshifting and compressing the pulse back to its original duration and shape. Notably, no chirping occurs here, given the uniformity of the space-time structure [37, 38].
@@
This Doppler-driven compansion mechanism sets DoPA apart from CPA. As the pulse propagates through the opening wedge, Doppler-induced downshifts gradually shift its spectral components toward lower frequencies, effectively narrowing the spectral width (Sec. III). This enables DoPA to operate with a much narrower amplifier bandwidth than CPA, thereby mitigating gain narrowing and preserving spectral integrity after amplification. Additionally, DoPA offers the potential for a compact implementation (Sec. V), which could be particularly advantageous in applications requiring high-power ultrashort pulses where system size and complexity are critical constraints.
@@ III System Analysis
Figure 2 shows a practical implementation of the double space-time wedge structure of the DoPA system (Fig. 1b) with corresponding spectral transition diagrams. Unlike idealized space-time wedges [36], this implementation includes space-time apertures that allow the pulse to enter and exit the structure without spurious reflections. In the opening wedge configuration (Fig. 2a), the input aperture is positioned beyond the wedge apex, while the output aperture is before the wedge basis. In the closing wedge configuration (Fig. 2b), the input and output apertures correspond to the time-reversed versions of their opening-wedge counterparts. Once inside the structure, the incident pulse—initially at $X_{0}=\left(z_{0},ct_{0}\right)$—undergoes multiple space-time reflections and related spectral transitions at distinct space-time scattering points $X_{1},X_{2},\ldots$, until its final reflection at $X_{p}$, after which it exits the structure. Both the opening and closing wedges move with a constant normalized modulation velocity, $\beta=v_{\text{m}}/c$, where $10<\beta<$ 1, which governs the Doppler shifts and temporal dynamics of the pulse as it propagates through the structure.
@@
To ensure full recovery of the original pulse shape and duration, the opening and closing wedges must be paired with precisely opposite space-time trajectories, i.e., the same velocity magnitude, and an identical number of reflections. This ensures the output retains the same duration and shape as the incident pulse. This design is analogous to chirp filters in CPA, where the phase modulation in the second filter compensates for the chirp introduced by the first (Fig. 1a). Likewise, in DoPA, the opening wedge modifies the pulse in a controlled manner, while the closing wedge precisely reverses these modifications.
@@
Figure 2: Implementation of the space-time wedges in the DoPA system of Fig. 1b with space-time apertures (top) and corresponding spectral transition diagrams (bottom). The wedges are the light-gray triangular areas, representing propagation media of unit refractive index ($n=1$), that compand pulses via Doppler shifting [Eq. (1)] by a compansion factor $a$ [Eq. (2)]. The two dark-gray areas sandwiching the wedges act as perfect electric conductors ($\sigma=\infty$). The modulation interfaces travel with normalized velocity $\beta=v_{\text{m}}/c$. The direct space-time points are denoted by $X_{p}=\left(z_{p},ct_{p}\right)$, while the inverse space-time points are denoted by $\tilde{X}_{p}=\left(k_{p},\omega_{p}/c\right)=R^{p}\tilde{X}_{0}$, where $R=\left(1\mp\beta\right)/\left(1\pm\beta\right)$ (Sec. A). (a) Opening wedge, expanding pulses. (b) Closing wedge, compressing pulses.
@@
When the incident pulse impinges on the moving interface, at $X_{1}$, it undergoes the Doppler frequency shift of a moving interface with constant velocity, $\beta$, viz., $\omega_{1}=R\omega_{0}$, where $R=\left(1\mp\beta\right)/\left(1\pm\beta\right)$ and $\omega_{0}$ is the center frequency of the incident wave [34] (Fig. 2). The top sign corresponds to an interface that is comoving with respect to the incident wave, resulting in $R<1$ and thus a decrease in frequency. The bottom sign corresponds to an interface that is contramoving with respect to the incident wave, resulting in $R>1$ and thus an increase in frequency.
@@
Within the wedge structure (Fig. 2), the pulse experiences multiple space-time reflections, each introducing its own Doppler shift. However, due the uniformity of the structure, the pulse undergoes the same Doppler shift at each scattering event. Therefore, the Doppler-shifted frequency, $\omega_{p}$, after $p$ space-time reflections is obtained by differentiating the phase argument of the pulse with respect to time and is given by (Sec. A.2):
@@
For an opening wedge (Fig. 2a), the wave encounters only comoving interfaces ($R<1)$ and thus the frequency decreases with each reflection. For a closing wedge (Fig. 2b), the pulse encounters only contramoving interfaces ($R>1$) and thus the frequency increases as the wave propagates through the structure.

### v3

@@ Abstract
The ability to amplify ultrashort pulses has revolutionized modern laser science, driving advances in various fields such as ultrafast optics and spectroscopy. A pivotal development in this field is chirped pulse amplification (CPA), which stretches, amplifies and recompresses ultrashort optical pulses using dispersive elements to overcome amplification limits. However, CPA faces limitations due to gain narrowing, restricting the final pulse duration. Here, we propose Doppler pulse amplification (DoPA), a novel approach for amplifying ultrashort pulses. While DoPA shares similarities with CPA in that it also stretches, amplifies and recompresses pulses, it differs in how it achieves this temporal compansion. Unlike CPA, DoPA exploits Doppler shifts induced by space-time modulated interfaces through a space-time wedge implementation without chirping. We show that DoPA dynamically shifts the pulse spectrum, effectively mitigating the gain narrowing issue of CPA. Additionally, we show that DoPA enables more compact amplification systems via a space-time Fresnel implementation. This approach may pave the way for more efficient, high-intensity laser systems and expand the potential for applications in both laboratory research and practical environments.
@@ I Introduction
Pulse amplification techniques have continuously evolved to meet the growing demands of various fields, from telecommunications and medical applications to scientific research. Some applications, such as communication [1, 2] and radar systems [3, 4], require only moderate gain amplification. However, others—including laser light-matter interactions [5, 6], particle acceleration [7, 8], ultrafast optics [6, 9] and precision eye surgery [10, 11]—demand advanced techniques to amplify ultrashort pulses to very high peak powers. In these cases, directly amplifying ultrashort pulses exceed the saturation level of the amplifier, potentially damaging optical components and introducing nonlinear effects such as self-phase modulation [12, 13]. To address this, two main approaches have been developed. The first, chirped pulse amplification (CPA) [14, 15, 16, 17, 18], uses dispersive elements, such as diffraction gratings, to temporally stretch the pulse, reducing its peak power and spreading its energy over time. This enables greater amplification before the pulse is recompressed to its original shape and duration, now with significantly higher peak power. Developed by Strickland and Mourou, CPA revolutionized laser science and earned them the 2018 Nobel Prize in Physics [14, 15]. The second approach, divided-pulse amplification (DPA) [19, 20, 21, 22, 23, 24, 25], splits the input pulse into multiple identical lower-power copies using a sequence of birefringent crystals. These copies are individually amplified and then recombined, resulting in a pulse with significantly higher peak power.
@@
Despite their success, both approaches come with intrinsic limitations that restrict their amplification performance. On one hand, CPA suffers from gain narrowing, a spectral bottleneck that arises when the spectral width of the pulse exceeds the bandwidth of the gain amplifier, leading to temporal broadening and pulse distortion in the output [26, 27, 18]. Additionally, CPA is constrained by its inability to accommodate the large diffraction grating sizes required for relatively long pulses  [^1] For a fixed amount of group velocity dispersion and a fixed stretching factor, the size of the diffraction grating scales with the square of the input pulse duration [9]—for instance, increasing the pulse duration from 101010 fs to 101010 ps requires enlarging the grating length by a factor of $10^{6}$—which renders diffraction grating technology impractical [21].. On the other hand, DPA is limited by the number of pulse divisions, restricting the reduction in peak power of the individual copies. Consequently, amplification is constrained, as the copies quickly reaches the saturation level of the amplifier [29, 30, 31].
@@
We propose a novel method for amplifying ultrashort pulses, called Doppler pulse amplification (DoPA). DoPA shares the same top-level structure with CPA—stretching, amplifying and compressing—but instead of using dispersive elements, it relies on Doppler frequency shifts [32] induced by moving modulated interfaces [33, 34, 35, 36] to control the pulse compansion (expansion and compression), without any chirping. It mitigates the gain narrowing problem by dynamically alternating the pulse spectrum, making it suitable for the amplification of ultrashort pulses, and may lead to compact systems.
@@
The paper is structured as follows. Section II introduces DoPA in parallel to CPA. Section III analyzes the DoPA system using a space-time wedge implementation. Section IV demonstrates this analysis. Section V discusses how the device size can be reduced by replacing the space-time wedge with a space-time Fresnel profile. Section VI outlines potential experimental realizations of the DoPA system. Finally, Sec. VII concludes the paper.
@@
Figure 1: General concept of (a) chirped pulse amplification (CPA) and (b) proposed Doppler pulse amplification (DoPA). In both CPA and DoPA, the input pulse is first stretched out in time to reduce its peak power from $P_{\text{in}}$ to $P^{\prime}_{\text{in}}=P_{\text{in}}/a$, where $a>1$ is the compansion factor. This temporal stretching—achieved via a chirp filter in CPA and an opening space-time wedge in DoPA—lowers the peak power enough to allow amplification up to the saturation threshold, $P_{\text{th}}$, of the amplifier. With gain, $G$, the amplified stretched pulse reaches $P^{\prime}_{\text{out}}=GP^{\prime}_{\text{in}}=P_{\text{th}}$. The pulse is then recompressed—using a matched chirp filter (CPA) and a matched closing space-time wedge (DoPA)—restoring its duration and boosting the peak power to $P_{\text{out}}=aP^{\prime}_{\text{out}}=GP_{\text{in}}$. Unlike CPA, where the spectral width remains constant through the chirp filter ($\Delta\omega_{\text{in}}^{\prime}=\Delta\omega_{\text{in}}$[9] and therefore, requires a broad amplifier bandwidth, DoPA dynamically compresses the spectrum during the stretching stage ($\Delta\omega_{\text{in}}^{\prime}\ll\Delta\omega_{\text{in}}$[36], enabling amplification with a much narrower-band amplifier.
@@ II Concept
Although based on a fundamentally different mechanism, DoPA shares the principle of temporal compansion with CPA. Therefore, it makes sense to introduce DoPA alongside CPA. Figure 1 depicts the two approaches, with Fig. 1a and 1b corresponding to CPA and DoPA, respectively.
@@
A CPA system, shown in Fig. 1a, consists of three main stages [14]. First, the input pulse—with peak power $P_{\text{in}}$—passes through a chirp filter characterized by a phase modulation $\varphi{\left[\omega\right]}$, whose group velocity dispersion temporally expands and chirps the pulse [9]. This causes different frequency components to experience different group delays, resulting in temporal stretching of the pulse. This chirping process reduces the peak power from $P_{\text{in}}$ to $P_{\text{in}}/a$, where $a>1$ is the compansion factor while spreading its energy over a longer duration. Second, the stretched pulse is amplified by a gain amplifier, e.g., Ti:sapphire or Nd:glass, up to its saturation threshold, $P_{\text{th}}$, to mitigate nonlinear effects and prevent damage to the optical components, resulting in an intermediate peak power $P^{\prime}_{\text{out}}=GP^{\prime}_{\text{in}}$, where $G$ is the amplifier gain. Finally, a second chirp filter, with opposite dispersion characteristics, compresses the pulse back to its original duration, eliminating the chirp introduced in the first step and significantly increasing the output peak power to $P_{\text{out}}=aP_{\text{out}}^{\prime}=GP_{\text{in}}$.
@@
Since the input pulse is ultrashort, it has a broad spectral width that remains unchanged after passing through the chirp filter [9]. This imposes strict constraints on the amplifier, whose limited bandwidth may amplify only a portion of the spectrum, leading to distortions that undermine the original goal of the system. This phenomenon, known as gain narrowing, is one of the principal limitations of CPA. In the figure, it is avoided by employing a very broadband amplifier, but it would arise if the narrower-band amplifier corresponding to the dotted curve in Fig. 1b were used [26, 27, 18].
@@
While DoPA also involves sequential pulse expansion, amplification and recompression, as illustrated in Fig. 1b, it achieves these transformations through Doppler frequency shifts induced by moving interfaces, instead of chirp filters. DoPA employs two paired space-time wedges, recently introduced in [36]. The wedge edges are assumed to act as perfect electric conductor moving interfaces, ensuring that the pulse remains confined within the structure. This configuration allows the pulse to bounce back and forth between two moving interfaces and depending on the wedge type—–either opening or closing—–the temporal envelope of the pulse either expands or compresses upon each interaction with the interfaces. In the case of an opening wedge, the pulse interacts only with comoving interfaces, leading to Doppler downshifting, which stretches the pulse duration by a factor $a$ and reduces its peak power to $P_{\text{in}}/a$, facilitating safe amplification. Conversely, for a closing wedge, the wave encounters only contramoving interfaces, resulting in Doppler upshifting and compressing the pulse back to its original duration and shape. Notably, no chirping occurs here, given the uniformity of the space-time structure [37, 38].
@@
This Doppler-driven compansion mechanism causes progressive spectral narrowing as the pulse propagates through the opening wedge (Sec. III), in contrast to CPA. This dynamic narrowing enables the use of much narrower-band amplifiers without incurring gain narrowing and preserving spectral integrity after amplification. Additionally, DoPA offers the potential for a compact implementation (Sec. V), which could be particularly advantageous in applications requiring high-power pulses where system size and complexity are critical constraints.
@@ III System Analysis
Figure 2 shows a practical implementation of the double space-time wedge structure of the DoPA system (Fig. 1b) with corresponding spectral transition diagrams. Unlike idealized space-time wedges [36], this implementation includes space-time apertures that allow the pulse to enter and exit the structure without spurious reflections. In the opening wedge configuration (Fig. 2a), the input aperture is positioned beyond the wedge apex, while the output aperture is before the wedge basis. In the closing wedge configuration (Fig. 2b), the input and output apertures correspond to the time-reversed versions of their opening-wedge counterparts. Once inside the structure, the incident pulse—initially at $X_{0}=\left(z_{0},ct_{0}\right)$—undergoes multiple space-time reflections and related spectral transitions at distinct space-time scattering points $X_{1},X_{2},\ldots$, until its final reflection at $X_{p}$, after which it exits the structure. Both the opening and closing wedges move with a constant normalized modulation velocity, $\beta=v_{\text{m}}/c$, where $0<\beta<1$, which governs the Doppler shifts and temporal dynamics of the pulse as it propagates through the structure.
@@
To ensure full recovery of the original pulse shape and duration, the opening and closing wedges must be paired with precisely opposite space-time trajectories, i.e., the same velocity magnitude, and an identical number of reflections. This design is analogous to chirp filters in CPA, where the phase modulation in the second filter compensates for the chirp introduced by the first (Fig. 1a). Likewise, in DoPA, the opening wedge modifies the pulse in a controlled manner, while the closing wedge precisely reverses these modifications.
@@
Figure 2: Implementation of the space-time wedges in the DoPA system of Fig. 1b with space-time apertures (top) and corresponding spectral transition diagrams (bottom). The wedges are the light-gray triangular areas, representing propagation media of unit refractive index ($n=1$), that compand pulses via Doppler shifting [Eq. (1)] by a compansion factor $a$ [Eq. (2)]. The two dark-gray areas sandwiching the wedges act as perfect electric conductors. The modulation interfaces travel with normalized velocity $\beta=v_{\text{m}}/c$. The direct space-time points are denoted by $X_{p}=\left(z_{p},ct_{p}\right)$, while the inverse space-time points are denoted by $\tilde{X}_{p}=\left(k_{p},\omega_{p}/c\right)$. (a) Opening wedge: pulses broaden temporally and compress spectrally. (b) Closing wedge: pulses compress temporally and broaden spectrally.
@@
When the incident pulse impinges on the moving interface, at $X_{1}$, it undergoes the Doppler frequency shift of a moving interface with constant velocity, $\beta$, viz., $\omega_{1}=R\omega_{0}$, where $R=\left(1\mp\beta\right)/\left(1\pm\beta\right)$ and $\omega_{0}$ is the center frequency of the incident wave [34] (Fig. 2). The top sign corresponds to an interface that is comoving with respect to the incident wave, resulting in $R<1$ and thus a decrease in frequency. The bottom sign corresponds to an interface that is contramoving with respect to the incident wave, resulting in $R>1$ and thus an increase in frequency. Because this scaling applies uniformly to the entire spectrum, both the center frequency, $\omega_{1}$, and the spectral width, $\Delta\omega_{1}$, are multiplied by the same factor, $R$, resulting in spectral compression for comoving interfaces ($R<1$) and spectral broadening for contramoving interfaces ($R>1$).
@@
Within the wedge structure (Fig. 2), the pulse experiences multiple space-time reflections, each introducing its own Doppler shift. However, due the uniformity of the structure, the pulse undergoes the same Doppler shift at each scattering event. Therefore, the Doppler-shifted frequency, $\omega_{p}$, after $p$ space-time reflections is simply (Sec. A.2)
@@
For an opening wedge (Fig. 2a), the wave encounters only comoving interfaces ($R<1)$ and thus both the frequency and spectral width decrease with each reflection. For a closing wedge (Fig. 2b), the pulse encounters only contramoving interfaces ($R>1$) and thus both the frequency and spectral width increase as the wave propagates through the structure.